OpenStudy (anonymous):

I cannot find where I messed up in this problem. Find the exact arc length of y=sqrt(x-x^2)+sin^-1(sqrt(x))

OpenStudy (anonymous):

from 0 to 1 i take it

OpenStudy (anonymous):

you need \[\int_0^1\sqrt{1+f'(x)^2}dx\] if i recall correctly, so first you need \[f'(x)=\frac{\sqrt{x-x^2}}{x}\]

OpenStudy (anonymous):

and so \[f'(x)^2=\frac{x-x^2}{x^2}=\frac{1}{x}-x\]

OpenStudy (anonymous):

add 1 and get \[\frac{-x^2+x+1}{x}\]

OpenStudy (anonymous):

damn that was wrong \[f'(x)^2=\frac{1}{x}-1\] so \[1+f'(x)^2=\frac{1}{x}\]

OpenStudy (anonymous):

now you have \[\int_0^1\sqrt{\frac{1}{x}}dx\] which should be ok right?

OpenStudy (anonymous):

no, there were not any intergration limits

OpenStudy (anonymous):

It wanted the exact value

OpenStudy (anonymous):

yea I got that

OpenStudy (anonymous):

but when I put in 2sqrt(x) as my answer, it said it was wrong

OpenStudy (anonymous):

well the funciton is only defined on the interval [0,1] so i assumed those were the limits of integration

OpenStudy (anonymous):

oh okay, I just assumed it meant no limits

OpenStudy (anonymous):

integral is 2

OpenStudy (anonymous):

i guess when it said "exact arc length" it meant the total arc length

OpenStudy (anonymous):

Haha, yay I got it. Thank for your help!!! Yea, I did not exactly understand that part. But it is all good now.

OpenStudy (anonymous):

great! yw